Nickblade10:

What is the simplified form of the expression the quantity x squared plus 10x plus 21 end quantity divided by 4 times the quantity x plus 3 end quantity? the quantity x minus 7 end quantity divided by 4 the quantity x plus 7 end quantity divided by 4 the quantity x plus 7 end quantity divided by quantity x plus 3 end quantity the quantity x minus 7 end quantity divided by quantity x plus 3 end quantity

Laylalyssa:

screenshot

Nickblade10:

working on it

Nickblade10:

Laylalyssa:

okay so all you need to do is simplify it so starting with the denominator you have 4(x+3) you have to do 4 times x, and then 4 times 3 you end up with 4x+12 after factoring you get \[\frac{ (x+3)(x+7) }{ 4(x+3) }\] that simplifies to \[\frac{ 1 }{ 4 }x +\frac{ 7 }{4 }\] which is \[\frac{ x+7 }{4 }\]

Laylalyssa:

do you get it?

Nickblade10:

processing it

Nickblade10:

Ok, well, i wrote it down in my notes